# Stochastic dominance for super heavy-tailed random variables

### v1

@@ 3 Super heavy-tailed distributions and stochastic dominance
If a random variable $X$ is super heavy-tailed, then it satisfies property (2). If $X$ is strictly super heavy-tailed, then $X<_{\rm st}\sum_{i=1}^{n}\theta_{i}X_{i}$.
@@
Let $X\sim F$ and $\bar{\theta}\in\Delta_{n}$. We have, for all $x>0$,
$\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}\leq x\right)$ $\leq\mathbb{P}(\theta_{1}X_{1}\leq x,\dots,\theta_{n}X_{n}\leq x)=\prod_{i=1}^{n}F\left(\frac{x}{\theta_{i}}\right)=\prod_{i=1}^{n}\exp\left(-h_{F}\left(\frac{\theta_{i}}{x}\right)\right)$
@@
An immediate consequence of Theorem 1 and Proposition 1 (i) is that super heavy-tailed distributions have infinite mean.
3.1 Properties of super heavy-tailed distributions
In this section, we study some properties of super heavy-tailed random variables. In particular, these properties will be used to demonstrate that many well-known infinite-mean random variables are super heavy-tailed and, thanks to further sufficient conditions for belonging to this class, one can construct many more super heavy-tailed random variables.
Proposition 2.
Let $X\sim F$ be a super heavy-tailed random variable. The following statements hold.
@@ 3.1 Properties of super heavy-tailed distributions
    $F$ is continuous on $[0,\infty)$.
  2. (ii)
    $F^{\beta}$, $\beta>0$, is super heavy-tailed.
  3. (iii)
    If, in addition, a random variable $Y\sim G$ is super heavy-tailed and independent of $X$, then $\max\{X,Y\}$ is also super heavy-tailed. In terms of distribution functions, if $F$ and $G$ are both super heavy-tailed, then so is $FG$.
  4. (iv)
    For a non-decreasing, convex, and non-constant function $f:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ with $f(0)=0$, $f(X)$ is super heavy-tailed.
Proof.
  1. (i)
    As $h_{F}$ is sub-additive and increasing, and $\lim_{x\downarrow 0}h_{F}(x)=0$, $h_{F}$ is continuous on $(0,\infty)$, and so is $F$ (see Remark 1 of Matkowski and Świątkowski, (1993)). The desired result is due to the right-continuity of $F$.
@@
    $\left(f^{-1+}\left(\frac{ab}{a+b}\right)\right)^{-1}\leq\left(f^{-1+}\left(a\right)\right)^{-1}+\left(f^{-1+}\left(b\right)\right)^{-1}.$ (3)
    Denote by $F$ and $G$ the distribution functions of $X$ and $f(X)$, respectively. Then $G(x)=\mathbb{P}(f(X)\leq x)=\mathbb{P}(X\leq f^{-1+}(x))=F(f^{-1+}(x))$ for $x\geq 0$. By letting $g(x)=1/f^{-1+}(1/x)$ for $x>0$, we write $h_{G}=h_{F}\circ g$. By inequality (3), for any $x,y>0$,
@@
    Therefore, $g$ is sub-additive. As $h_{F}$ is sub-additive and non-decreasing, it is clear that $h_{G}=h_{F}\circ g$ is sub-additive and we have the desired result. ∎
Proposition 3.
Let $\bar{\theta}\in\Delta_{n}$. If distribution functions $F_{1},\dots,F_{n}$ are super heavy-tailed and $F_{1}\leq_{\rm st}\dots\leq_{\rm st}F_{n}$, then $\sum_{i=1}^{n}\theta_{i}F_{i}$ is also super heavy-tailed.
Proof.

### v3

@@ 3.3 Main result
If a random variable $X\in\mathcal{H}$ and random variables $X_{1},\dots,X_{n}$ are NLOD with marginal laws equal to $X$, then for $\bar{\theta}\in\Delta_{n}^{+}$,
$X\leq_{\rm st}\theta_{1}X_{1}+\dots+\theta_{n}X_{n}.$ (2)
If $X\in\mathcal{H}_{s}$, then $X<_{\rm st}\sum_{i=1}^{n}\theta_{i}X_{i}$.
@@
Let $X\sim F$ and $\bar{\theta}\in\Delta_{n}^{+}$. We have, for all $x>0$,
@@
An immediate consequence of Theorem 1 and Proposition 1 (i) is that all distributions in $\mathcal{H}$ have infinite mean.
Remark 2 (Value-at-Risk).
One regulatory risk measure in insurance and finance is Value-at-Risk (VaR). For a random variable $X\sim F$ and $p\in(0,1)$, VaR is defined as $\mathrm{VaR}_{p}(X)=F^{-1}(p)$. For two random variables $X$ and $Y$, it is well known that $X\leq_{\rm st}Y$ if and only if $\mathrm{VaR}_{p}(X)\leq\mathrm{VaR}_{p}(Y)$ for all $p\in(0,1)$. Note that VaR is comonotonic-additive; a risk measure $\rho$ is comonotonic-additive if $\rho(Y+Z)=\rho(Y)+\rho(Z)$ for comonotonic random variables $Y$ and $Z$ [^1] Random variables $Y$ and $Z$ are comonotonic if there exists a random variable $U$ and two increasing functions $f$ and $g$ such that $Y=f(U)$ and $Z=g(U)$ almost surely. Then we have $\mathrm{VaR}_{p}(X)=\sum_{i=1}^{n}\mathrm{VaR}_{p}(\theta_{i}X_{i})$ for identically distributed random variables $X_{1},\dots,X_{n}$. By Theorem 1, superadditivity of VaR holds for $\bar{\theta}\in\Delta_{n}^{+}$ and identically distributed risks $X_{1},\dots,X_{n}\in\mathcal{H}$ that are NLOD: For all $p\in(0,1)$, $\mathrm{VaR}_{p}(\theta_{1}X_{1})+\dots+\mathrm{VaR}_{p}(\theta_{n}X_{n})\leq\mathrm{VaR}_{p}(\theta_{1}X_{1}+\dots+\theta_{n}X_{n}).$ More generally, the superadditivity property holds for any comonotonic-additive risk measure that is consistent with stochastic order.
Remark 3 (Heavy-tailed distributions).
A distribution $F$ is said to be heavy-tailed in the sense of Falk et al., (2011) with tail parameter $\alpha>0$, if $\overline{F}(x)=L(x)x^{-\alpha}$ where $L$ is a slowly varying function, that is, $L(tx)/L(x)\rightarrow 1$ as $x\rightarrow\infty$ for all $t>0$. For iid heavy-tailed random variables $X_{1},X_{2},\dots\sim F$, if there exist sequences of constants $\{a_{n}\}$ and $\{b_{n}\}$ where $b_{n}>0$ such that $(\max\{X_{1},\dots,X_{n}\}-a_{n})/b_{n}$ converges to the Fréchet distribution, $F$ is said to be in the maximum domain attraction of the Fréchet distribution. It is known in the Extreme Value Theory (Embrechts et al. (1997)) that a distribution is in the maximum domain of attraction of the Fréchet distribution if and only if the distribution is heavy-tailed. Note that for a heavy-tailed random variable $X$ with $\alpha\leq 1$, $\mathbb{E}(|X|)=\infty$. An interesting property of heavy-tailed risks with infinite mean is the asymptotic superadditivity of VaR: If $X_{1},\dots,X_{n}$ are iid and heavy-tailed with tail parameter $\alpha<1$,
$\lim_{p\rightarrow 1}\frac{\mathrm{VaR}_{p}(X_{1}+\dots+X_{n})}{\mathrm{VaR}_{p}(X_{1})+\dots+\mathrm{VaR}_{p}(X_{n})}>1.$
See, e.g., Example 3.1 of Embrechts et al. (2009) for the claim above. Heavy-tailed risks with infinite mean are not necessarily in $\mathcal{H}$ as the condition of $\mathcal{H}$ applies over the whole range of distributions whereas heavy-tailed distributions have power-law shapes only in their tail parts. On the other hand, risks in $\mathcal{H}$ are not necessarily heavy-tailed in the sense of Falk et al., (2011). For instance, the survival distributions of log-Pareto risks are slowly varying functions. Distributions with slowly varying tails are called super heavy-tailed.
Remark 4 (Convex order).
Besides stochastic order, another popular notion of stochastic dominance to compare risks is convex order. For two random variables $X$ and $Y$, $X$ is said to be smaller than $Y$ in convex order, denoted by $X\leq_{\rm cx}Y$, if $\mathbb{E}(u(X))\leq\mathbb{E}(u(Y))$ for all convex functions $u$ provided that the expectations exist. The interpretation of $X\leq_{\rm cx}Y$ is that $Y$ is more “spread-out” than $X$. If $X_{1},\dots,X_{n}$ are iid and have a finite mean, by Theorem 3.A.35 of Shaked and Shanthikumar (2007), for $\bar{\theta}\in\Delta_{n}^{+}$, $\sum_{i=1}^{n}\theta_{i}X_{i}\leq_{\rm cx}X_{1}$. Unlike Theorem 1, this leads to a diversification benefit. Note that $\leq_{\rm cx}$ is not suitable for the analysis of risks with infinite mean as the expectation of any increasing convex transform of these risks is infinity.
Remark 5 (Positive dependence).
One may expect positive dependence to make larger values of the sum in (2) more likely and thus the sum more likely to stochastically dominate a single random variable. We believe that this intuition does not hold due to the very heavy tails of the random variables under consideration. It is known, for instance, that very large values of the sum of iid random variables with heavy tails are likely caused by a single random variable taking a large value, while other random variables are moderate. If random variables are positively dependent and some of them do not take large values, it makes others more likely to take moderate values too, hence positive dependence may hinder large values; see Alink et al., (2004) and Mainik and Rüschendorf, (2010) for such observations in some asymptotic senses. These phenomena can also be seen from the deadly risks considered by Müller, (2024): For all $i\in[n]$, $\mathbb{P}(X_{i}=0)=1-p$ and $\mathbb{P}(X_{i}=\infty)=p$ where $p>0$. For $\overline{\theta}\in\Delta_{n}^{+}$, it is clear that $\sum_{i=1}^{n}\theta_{i}X_{i}=\infty$ as long as one of $X_{1},\dots,X_{n}$ is $\infty$. If $X_{1},\dots,X_{n}$ are positively lower orthant dependent (PLOD), that is $\mathbb{P}(X_{1}\leq x_{1},\dots,X_{n}\leq x_{n})\geq\prod_{i=1}^{n}\mathbb{P}(X_{i}\leq x_{i})$ for all $x_{1},\dots,x_{n}\in\mathbb{R}$, then
$\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}=\infty\right)$ $=1-\mathbb{P}\left(X_{1}=\dots=X_{n}=0\right),$
$=1-\mathbb{P}\left(X_{1}\leq 0,\dots,X_{n}\leq 0\right)\leq 1-\prod_{i=1}^{n}\mathbb{P}(X_{i}=0).$
Hence, $\sum_{i=1}^{n}\theta_{i}X_{i}$ is stochastically smaller when $X_{1},\dots,X_{n}$ are PLOD compared to the case when $X_{1},\dots,X_{n}$ are independent. The situation is reversed for NLOD random variables. However, (SD) still holds for PLOD risks $X_{1},\dots,X_{n}$ as
$\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}=\infty\right)=1-\mathbb{P}\left(X_{1}=\dots=X_{n}=0\right)\geq 1-\mathbb{P}(X_{1}=0)=\mathbb{P}(X_{1}=\infty).$
In Chen et al., (2025), (SD) is shown to hold for infinite-mean Pareto random variables that are positively dependent via some specific Clayton copula.
4 Weighted sums of non-identically distributed risks
In the previous section, property (SD) is studied for risks with the same marginal distribution. We now look at the case when risks are not necessarily identically distributed. Given non-identically distributed random variables $X_{1},\dots,X_{n}$ and any $\bar{\theta}\in\Delta_{n}$, the question is to study for which random variable $X$ the following property holds
$X\leq_{\rm st}\theta_{1}X_{1}+\dots+\theta_{n}X_{n}.$ (3)
To study this problem, we introduce the class of super-Fréchet distributions defined below.
Definition 4.
A random variable $X$ is said to be super-Fréchet (or has a super-Fréchet distribution) if $X$ and $f(Y)$ have the same distribution, where $Y\sim$ Fréchet$(1)$ and $f$ is a strictly increasing and convex function with $f(0)=0$.
As convex transforms make the tail of random variables heavier, super-Fréchet distributions are more heavy-tailed than Fréchet$(1)$ distribution, and thus the name. It is easy to check that a random variable $X$ with $\operatorname*{ess\text{-}inf}X=0$ is super-Fréchet if and only if the function $g:x\mapsto 1/(-\log\mathbb{P}(X\leq x))$ is strictly increasing and concave on $(0,\infty)$ with $\lim_{x\downarrow 0}g(x)=0$.
As Fréchet$(1)$ distribution is in $\mathcal{H}$, by Proposition 1 (iv), Super-Fréchet distributions are in $\mathcal{H}$. On the other hand, not all distributions in $\mathcal{H}$ are Super-Fréchet, which can be seen from the following example.
Example 3.
For $c>0$, define a distribution function on $(0,\infty]$:
$F(x)=\exp(-c\lceil 1/x\rceil),\mbox{~{}~{}for~{}~{}}x\in(0,\infty),$
and $F(\infty)=1$. Then $h_{F}(x)=c\lceil x\rceil$, $x\in(0,\infty)$, is subadditive, and thus $F\in\mathcal{H}$. However, since $F$ is not a continuous distribution, it is not super-Fréchet. The distributions in this example are the so-called inverse-geometric distributions, also considered in Example 2.7 of Arab et al., (2024).
Fréchet distributions with infinite mean, as well as many other distributions in the following example, are super-Fréchet.
Example 4.
Pareto, Burr, paralogistic, and log-logistic random variables, all with infinite mean, are super-Fréchet distributions. Since all these random variables can be obtained by applying strictly increasing and convex transforms to Pareto$(1)$ random variables (see Appendix C), it suffices to show that a Pareto$(1)$ random variable is super-Fréchet. Write the Pareto$(1)$ distribution as $F(x)=1-1/(x+1)=\exp(-1/g(x))$, $x>0$, where $g(x)=1/\log(1+1/x)$. It is clear that $g$ is strictly increasing and $\lim_{x\downarrow 0}g(x)=0$. We show $g$ is concave on $(0,\infty)$. We have
The generalized 00-mean function is the weighted geometric mean, that is, $M^{\mathbf{w}}_{0}(u_{1},\dots,u_{n})=\prod_{i=1}^{n}u_{i}^{w_{i}},$ which is also the limit of $M^{\mathbf{w}}_{r}$ as $r\to 0$. A generalized mean of distribution functions is a distribution function. In particular, if $r=1$, it leads to a distribution mixture model, i.e., if $X\sim M^{\mathbf{w}}_{1}(F_{1},\dots,F_{n})$, $X$ has the same distribution as $\sum_{i=1}^{n}X_{i}\mathds{1}_{A_{i}}$ where $A_{1},\dots,A_{n}$ are mutually exclusive, independent of $X_{1},\dots,X_{n}$, and $\mathbb{P}(A_{i})=w_{i}$ for all $i\in[n]$.
Theorem 2.
If $X_{1},\dots,X_{n}$ are super-Fréchet and NLOD with $X_{i}\sim F_{i}$, $i\in[n]$, and $X\sim M^{\bar{\theta}}_{r}(F_{1},\dots,F_{n})$ for some $r\geq 0$, then for $\bar{\theta}\in\Delta_{n}^{+}$,
$X\leq_{\rm st}\theta_{1}X_{1}+\dots+\theta_{n}X_{n}.$
Proof.
$=M^{\bar{\theta}}_{0}(F_{1}(x),\dots,F_{n}(x))\leq M^{\bar{\theta}}_{r}(F_{1}(x),\dots,F_{n}(x))=\mathbb{P}(X\leq x).$
5 Comparison with existing results
In this section, we compare our results with the literature. We first consider the case when $X_{1},\dots,X_{n}$ in (SD) are iid. In Arab et al., (2024), it is shown that (SD) holds for non-negative random variables that are InvSub; a random variable $X\sim F$ and its distribution is called InvSub if $1-F(1/x)$ is subadditive. The class of InvSub distributions is larger than $\mathcal{H}$ as $h_{F}=-\log F(1/x)$ is subadditive implies that $1-F(1/x)$ is subadditive. Müller, (2024) showed that (SD) holds for super-Cauchy random variables; a random variable $X\sim F$ and its distribution is called super-Cauchy if $F^{-1}(G(x))$ is convex where $G$ is the standard Cauchy distribution function. Super-Cauchy distributions are continuous and can take positive values on the entire real line but they do not contain $\mathcal{H}$ as $\mathcal{H}$ includes non-continuous distributions (see Example 3). The proofs in both Arab et al., (2024) and Müller, (2024) are short and elegant.
As our results cover the case of negatively dependent risks, for the rest of this section, we will focus on the comparison of our results with Chen et al. (2024a); to our best knowledge, Chen et al. (2024a) is the only other paper that deals with negatively dependent risks. Chen et al. (2024a) showed that
(SD) holds for super-Pareto risks X1,,Xn that are weakly negatively associated.italic-(SDitalic-) $\mbox{~{}holds for super-Pareto risks $X_{1},\dots,X_{n}$ that are weakly negatively associated.}$. (5)
For ease of comparison, definitions of super-Pareto distributions and weak negative association are given in a slightly different form from Chen et al. (2024a) below.
Definition 5.
A random variable $X$ and its distribution is super-Pareto if $X$ and $f(Y)$ have the same distribution for some non-decreasing, convex, and non-constant function $f$ with $f(0)=0$ and $Y\sim\mathrm{Pareto}(1)$.
Definition 6.
A set $S\subseteq\mathbb{R}^{k}$, $k\in\mathbb{N}$ is decreasing if $\mathbf{x}\in S$ implies $\mathbf{y}\in S$ for all $\mathbf{y}\leq\mathbf{x}$. Random variables $X_{1},\dots,X_{n}$ are weakly negatively associated if for any $i\in[n]$, decreasing set $S\subseteq\mathbb{R}^{n-1}$, and $x\in\mathbb{R}$ with $\mathbb{P}(X_{i}\leq x)>0$,
$\mathbb{P}(\mathbf{X}_{-i}\in S,X_{i}\leq x)\leq\mathbb{P}(\mathbf{X}_{-i}\in S)\mathbb{P}(X_{i}\leq x).$
where $\mathbf{X}_{-i}=(X_{1},\dots,X_{i-1},X_{i+1},\dots,X_{n})$.
Lemma 1.
If random variables $X_{1},\dots,X_{n}$ are super-Pareto and weakly negatively associated, then $X_{1},\dots,X_{n}\in\mathcal{H}$ and they are NLOD.
Proof.
As Pareto$(1)$ risks are in $\mathcal{H}$ (see Example 2), by Proposition 2 (iv), super-Pareto risks are in $\mathcal{H}$. Since $X_{1},\dots,X_{n}$ are weakly negatively associated, for any $(x_{1},\dots,x_{n})\in\mathbb{R}^{n}$,
$\mathbb{P}(X_{1}\leq x_{1},\dots,X_{n}\leq x_{n})\leq\mathbb{P}(X_{1}\leq x_{1},\dots,X_{n-1}\leq x_{n-1})\mathbb{P}(X_{n}\leq n)\leq\prod_{i=1}^{n}\mathbb{P}(X_{i}\leq x_{i}).$
Thus, $X_{1},\dots,X_{n}$ are NLOD. ∎
The above lemma shows that Theorem 1 implies (5), which is in Theorem 1 (i) of Chen et al. (2024a). We present below a corollary, which leads to a similar result as Theorem 1 (ii) of Chen et al. (2024a).
Corollary 1.
Suppose that a random variable $X\in\mathcal{H}$, random variables $X_{1},\dots,X_{n}$ are NLOD with marginal laws equal to $X$, and $\xi_{1},\dots,\xi_{n}$ are any positive random variables independent of $X,X_{1},\dots,X_{n}$ with $\sum_{i=1}^{n}\xi_{i}\leq 1$. If $\mathbb{P}(cX>t)\geq c\mathbb{P}(X>t)$ for all $c\in(0,1]$ and $t>0$, then for $x\geq 0$,
Proof.
By Theorem 1 and the independence between $\xi_{1},\dots,\xi_{n}$ and $X,X_{1},\dots,X_{n}$, we have
$\mathbb{P}\left(\sum_{i=1}^{n}\xi_{i}X_{i}>x\right)$ $=\mathbb{E}\left[\mathbb{P}\left(\sum_{i=1}^{n}\xi_{i}X_{i}>x|(\xi_{1},\dots,\xi_{n})\right)\right]$
$\geq\mathbb{E}\left[\mathbb{P}\left(\left(\sum_{i=1}^{n}\xi_{i}\right)X>x|(\xi_{1},\dots,\xi_{n})\right)\right]\geq\mathbb{E}\left(\sum_{i=1}^{n}\xi_{i}\right)\mathbb{P}\left(X>x\right).\qed$
For $\bar{\theta}\in\Delta_{n}$, let $A_{1},\dots,A_{n}$ be any events independent of $(X_{1},\dots,X_{n})$ and event $A$ be independent of $X$ satisfying $\mathbb{P}(A)=\sum_{i=1}^{n}\theta_{i}\mathbb{P}(A_{i})$. If $X_{1},\dots,X_{n}$ are financial losses, $A_{1},\dots,A_{n}$ can be interpreted as the triggering events for these losses. Let $\xi_{i}=\theta_{i}\mathds{1}_{A_{i}}$. By (6), for $x\geq 0$,
$\mathbb{P}\left(\sum_{i=1}^{n}\theta_{i}X_{i}\mathds{1}_{A_{i}}>x\right)\geq\mathbb{E}\left(\sum_{i=1}^{n}\theta_{i}\mathds{1}_{A_{i}}\right)\mathbb{P}(X>x)=\mathbb{P}(A)\mathbb{P}(X>x)=\mathbb{P}(X\mathds{1}_{A}>x),$
which is equivalent to
$X\mathds{1}_{A}\leq_{\rm st}\sum_{i=1}^{n}\theta_{i}X_{i}\mathds{1}_{A_{i}}.$ (7)
Theorem 1 (ii) of Chen et al. (2024a) showed (7) with different assumptions from Corollary 1; we refer readers to Chen et al. (2024a) for more details.
6 Conclusion
In this paper, we provide some sufficient conditions for property (SD) to hold. One can see that the property, while very strong, holds for a remarkably large class of distributions. We have also shown that it remains valid for non-identically distributed random variables.
We conclude with some open questions. First, we are interested in understanding how close our sufficient conditions for (SD) are to the optimal ones, i.e., we would like to understand what conditions are necessary for (SD).
Second, the definition of our class of heavy-tailed random variables seems to suggest that it is the distribution of $1/X$ that is of importance. We currently lack an intuitive explanation of this.
Finally, property (SD) raises the possibility that, for some random variables $X_{1},\dots,X_{n}$ and two vectors $\bar{\eta},\bar{\gamma}\in\mathbb{R}^{n}_{+}$,
where $\bar{\gamma}$ is smaller than $\bar{\eta}$ in majorization order; that is, $\sum^{n}_{i=1}\gamma_{i}=\sum^{n}_{i=1}\eta_{i}$ and $\sum^{k}_{i=1}\gamma_{(i)}\geq\sum^{k}_{i=1}\eta_{(i)}\,$ for $k\in[n-1]$ where $\gamma_{(i)}$ and $\eta_{(i)}$ represent the $i$th smallest order statistics of $\bar{\gamma}$ and $\bar{\eta}$. Clearly, (8) implies (SD). It is well known that (8) holds for iid stable random variables with infinite mean (see Ibragimov, (2005)) and it was recently shown to hold for iid Pareto random variables with infinite mean by Chen et al., (2025). It is of question whether (8) can hold for a larger class of distributions. Note that the methods used in the current paper do not appear to be useful to address (8) as we rely on the comparison of a sum with each of the summands. A more subtle approach to sums is required.
Appendices
Appendix A Proof of Proposition 2
@@
    As $h_{F}$ is subadditive and increasing, and $\lim_{x\downarrow 0}h_{F}(x)=0$, $h_{F}$ is continuous on $(0,\infty)$, and so is $F$ (see Remark 1 of Matkowski and Świątkowski, (1993)). The desired result is due to the right-continuity of $F$.
@@
    $\left(f^{-1+}\left(\frac{ab}{a+b}\right)\right)^{-1}\leq\left(f^{-1+}\left(a\right)\right)^{-1}+\left(f^{-1+}\left(b\right)\right)^{-1}.$ (A.1)
    Denote by $F$ and $G$ the distribution functions of $X$ and $f(X)$, respectively. Then $G(x)=\mathbb{P}(f(X)\leq x)=\mathbb{P}(X\leq f^{-1+}(x))=F(f^{-1+}(x))$ for $x\geq 0$. By letting $g(x)=1/f^{-1+}(1/x)$ for $x>0$, we write $h_{G}=h_{F}\circ g$. By inequality (A.1), for any $x,y>0$,
@@
    Therefore, $g$ is subadditive. As $h_{F}$ is subadditive and non-decreasing, it is clear that $h_{G}=h_{F}\circ g$ is subadditive and we have the desired result.
Appendix B Proof of Proposition 3